# TokenRec: Learning to Tokenize ID for LLM-based Generative Recommendations

### v2

@@ Abstract
There is a growing interest in utilizing large language models (LLMs) to advance next-generation Recommender Systems (RecSys), driven by their outstanding language understanding and reasoning capabilities. In this scenario, tokenizing users and items becomes essential for ensuring seamless alignment of LLMs with recommendations. While studies have made progress in representing users and items using textual contents or latent representations, challenges remain in capturing high-order collaborative knowledge into discrete tokens compatible with LLMs and generalizing to unseen users/items. To address these challenges, we propose a novel framework called TokenRec, which introduces an effective ID tokenization strategy and an efficient retrieval paradigm for LLM-based recommendations. Our tokenization strategy involves quantizing the masked user/item representations learned from collaborative filtering into discrete tokens, thus achieving smooth incorporation of high-order collaborative knowledge and generalizable tokenization of users and items for LLM-based RecSys. Meanwhile, our generative retrieval paradigm is designed to efficiently recommend top-K items for users, eliminating the need for the time-consuming auto-regressive decoding and beam search processes used by LLMs, thus significantly reducing inference time. Comprehensive experiments validate the effectiveness of the proposed methods, demonstrating that TokenRec outperforms competitive benchmarks, including both traditional recommender systems and emerging LLM-based recommender systems.
@@
Recommender Systems, Large Language Models, ID Tokenization, Vector Quantization, Collaborative Filtering.
@@ I Introduction
As a prominent branch in the data mining field, recommender systems (RecSys) serve as an indispensable and effective technique in addressing information overload problems and enriching user experience across diverse applications [1, 2], such as e-commerce, job search, and social media platforms. To provide personalized recommendations that accord with user preferences, one of the most representative techniques is collaborative filtering (CF), which aims to capture collaborative knowledge by modeling the history of user-item interactions [3]. For example, as the most classic CF method, Matrix Factorization (MF) [4] decomposes the use-item matrix into two low-rank matrices (i.e., user&item representations) to represent each user and item with a unique ID (namely ID indexing/tokenizing), and calculates matching scores to predict user behaviors and generate recommendations via the inner product between their representations. Due to the superior ability in representation learning on graph-structured data, Graph Neural Networks (GNNs) (e.g., LightGCN [5] and GTN [6]) have been recently employed to significantly enhance CF by capturing higher-order collaborative knowledge on user-item interaction graphs for recommendations in IDs manners. The main idea of the existing methods is to obtain learnable representations (i.e., token embeddings) for discrete user&item IDs (i.e., tokens).
@@
More recently, the rapid development of Large Language Models (LLMs) techniques (e.g., ChatGPT and LLaMA) has showcased notable milestones for revolutionizing natural language processing techniques [7, 8]. Technically, the unprecedented power of LLMs can be attributed to the scaling up of model/parameter size alongside a tremendous amount of training corpus. In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation abilities, along with remarkable generalization capability and in-context learning skills that facilitate LLMs to better generalize to unseen tasks and domains [8]. Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems [2]. For example, ChatRec [9] integrates ChatGPT into conversational recommendations, which enables users to deliver their explicit preferences in natural language. P5 [10] introduces an LLM-based recommendation framework that unifies diverse recommendation tasks by multi-task prompt-based pre-training.
@@ II-C3 $K$-way Encoder and Codebook
[…]
@@ II-D1 Tokenization&Prompts
Technically, Prompt 1 and Prompt 2 showcase the input prompts without and with item interactions as supporting information, respectively, in which $-\left\langle uk\text{-}\cdot\right\rangle$ denotes the out-of-vocabulary (OOV) tokens in the $k_{th}$ sub-codebook. For instance, $\left\langle u2\text{-}21\right\rangle$ represents the $21_{st}$ token in the second sub-codebook for the user_03. This also applies to tokenizing items in LLMs.
@@ II-D2 User Modeling via LLM
[…]
@@
In other words, the final encoded representation $\mathbf{h}_{i}$ can be considered as user $u_{i}$’s generative preferences of the next items for making personalized recommendations.
@@ II-D3 Generative Retrieval
[…]
@@
Given the scores of all items in the item base, the proposed TokenRec can easily retrieve top-$K$ items for users rather than directly generating tokens in natural language, thus offering advantages in terms of efficiency in inference and avoiding hallucinations. Furthermore, unseen items in the fine-tuning stage can be retrieved by only updating the item representations pool $\mathcal{V}$ instead of retraining the entire LLM model.
@@ II-E1 Training
Figure 3: The TokenRec’s efficiency and generalization capability for new users and items during the inference stage. Rather than retraining the entire model, which can be computationally expensive and time-consuming, only the GNN needs to be updated for learning representations for new users and items while keeping the MQ-Tokenizers and LLM backbone frozen. This significantly reduces the computational burden and allows faster adaptation to new users and items.
@@ II-E2 Inference
  • Generalizability to New Users and Items. The proposed architecture can provide robust ID tokenization for unseen users and items without fine-tuning the LLM4Rec component. As shown in Figure 3, when new users and items are added to the system, retraining is only required for the collaborative filtering component (e.g., MF and LightGCN) to capture collaborative knowledge in learning representations of these users and items (i.e., updating vector database). In contrast, the MQ-Tokenizers and LLM backbone can remain frozen and perform well, thanks to the masking and $K$-way encoder mechanisms in vector quantization. This eliminates the need for retraining LLM-related components and therefore saves huge computational resources: compared to the finetuning of LLMs, the training of GNNs is far more efficient [29]. Notably, the capability to adapt to new users and items is demonstrated in Section III-C and Table IV.
@@ III-A2 Baselines
Here, we compare our approach with four representative collaborative filtering methods (i.e., MF, NeuCF, LightGCN, GTN, and LTGNN), three widely-used sequential recommendation methods (i.e., SASRec, BERT4Rec, and S3Rec), and four state-of-the-art LLM-based recommendation methods(i.e., P5, CID, POD, and CoLLM).
  • Collaborative Filtering: MF [33] is the most classic CF method, while NeuCF [34] is the very first DNNs based collaborative filtering. LightGCN [5] and GTN [6] are representative CF based on GNNs techniques. LTGNN [29] is the most advanced GNN-based collaborative filtering method.
  • Sequential Recommendations: SASRec [35] is a self-attention based sequential recommendation model. BERT4Rec [36] is a bidirectional Transformer-based recommender trained with the BERT-style cloze task. S3Rec [37] is a representative sequential recommendation model trained by self-supervised learning.
@@
  • CID [38] is a non-trivial indexing approach that considers the co-occurrence matrix of items to design numeric IDs so that items co-occur in user-item interactions will have similar numeric IDs. To be consistent, we employ the P5 model as its LLM backbone.
@@
  • CoLLM [14] employs GNNs to provide continuous embeddings representing items and users for LLM-based recommendations. Notably, in our experiments, CoLLM’s binary classification output (i.e., whether or not a user likes an item) is reformatted to generate item IDs consistent with the output setting of top-$K$ recommendations.
@@ III-A3 Evaluation Metrics
In order to evaluate the quality of recommendation results, we adopt two widely used evaluation metrics: top-$K$ Hit Ratio (HR@$K$) and top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$[18, 29], in which higher values indicate better performance for recommendations. And the average metrics for all users in the test set are reported. In addition, we set the values of $K$ as 10, 20, and 30, among which 20 is the default value for ablation experiments.
@@ III-A4 Hyper-parameter Settings
Our proposed model is implemented based on Hugging Face and PyTorch. The codebook number $K$ (i.e., the subencoder), the token number $L$ at each subcodebook, and the ratio $\rho$ of our masking operation are search in the ranges of $\{1,2,3,4,5\}$, $\{128,256,512,1024\}$, and {0 to 1} in 0.1 increments, respectively. Moreover, during the fine-tuning process of our proposal LLM-based recommendation framework, the ratio of negative sampling $\lambda$ presented in Eq.(20) is fixed at 1:1, in which we randomly select an un-interacted item from the whole item base as the negative sample for each positive sample. The margin $\gamma$ in Eq.(20) is set to 0 to 0.2. We optimize the MQ-Tokenizers and the LLM backbone with AdamW [40] in a mini-batch manner. Note that the high-order collaborative representations for users and items are obtained from a representative collaborative filtering method, namely LightGCN [5]. For prompting, we design 11 templates for TokenRec: 10 of them are set to be seen prompts, and the remaining one is evaluated as an unseen prompt. For a fair comparison, we employ a widely-used lightweight LLM, i.e., T5-small [32], for TokenRec and all LLM-based baselines. The other default hyper-parameters for baseline methods are set as suggested by the corresponding papers.
@@ III-B Performance Comparison of Recommender Systems
  • As an earlier non-trivial indexing method to capture hidden knowledge in co-occurrence frequency, Collaborative Indexing (CID) outperforms Random Indexing (RID) and Sequential Indexing (SID) given the same setting of P5. These observations suggest the potential of integrating collaborative knowledge for item&user tokenization/indexing. However, the P5 variants and POD are inferior to the existing GNN-based collaborative filtering (i.e., LightGCN, GTN, and LTGNN), implying their inability to capture collaborative information using LLMs.
  • The GNN-based collaborative filtering methods perform relatively better than the traditional CF methods (i.e., MF and NCF) and representative sequential recommendation methods (i.e., BERT4Rec, SASRec, and S3Rec). The results demonstrate the effectiveness of GNNs in capturing collaborative signals via high-order connectivity.
@@ III-C Generalizability Evaluation
For this analysis, we consider the Beauty and LastFM datasets and exclude the 5% of users (ref as unseen users) with the least interaction history from the training data split to simulate newly added users. In this evaluation, the performance of TokenRec is compared to other LLM-based recommendation models, i.e., P5, POD, CID, and CoLLM. To ensure the absence of data leakage concerning unseen users, only the training split is utilized for establishing these LLM4Rec models, while it is allowed to update the vector database to provide collaborative representations of unseen users and items (as shown in Figure 3), which consumes far less computational resources than updating the LLM backbone of LLM-based RecSys. From the recommendation results presented in Table IV, several key observations can be made:
@@
  • In comparison, TokenRec outperforms all the compared methods for not only the training users but also the unseen users in both datasets. For instance, in the Amazon-Beauty dataset, the performance of TokenRec decreases by only 7% on average, demonstrating the strong generalization capability of TokenRec for newly added users. Such a superiority can be attributed to our MQ-Tokenizer for robust ID tokenization and the generative retrieval paradigm for flexible recommendation generation.

### v3

@@ Abstract
There is a growing interest in utilizing large language models (LLMs) to advance next-generation Recommender Systems (RecSys), driven by their outstanding language understanding and reasoning capabilities. In this scenario, tokenizing users and items becomes essential for ensuring seamless alignment of LLMs with recommendations. While studies have made progress in representing users and items using textual contents or latent representations, challenges remain in capturing high-order collaborative knowledge into discrete tokens compatible with LLMs and generalizing to unseen users/items. To address these challenges, we propose a novel framework called TokenRec, which introduces an effective ID tokenization strategy and an efficient retrieval paradigm for LLM-based recommendations. Our tokenization strategy involves quantizing the masked user/item representations learned from collaborative filtering into discrete tokens, thus achieving smooth incorporation of high-order collaborative knowledge and generalizable tokenization of users and items for LLM-based RecSys. Meanwhile, our generative retrieval paradigm is designed to efficiently recommend top-K items for users, eliminating the need for the time-consuming auto-regressive decoding and beam search processes used by LLMs, thus significantly reducing inference time. Comprehensive experiments validate the effectiveness of the proposed methods, demonstrating that TokenRec outperforms competitive benchmarks, including both traditional recommender systems and emerging LLM-based recommender systems. Codes and data are available at https://github.com/Quhaoh233/TokenRec.
@@
Recommender Systems, Large Language Models, ID Tokenization, Vector Quantization, and Collaborative Filtering.
@@ I Introduction
As a prominent branch in the data mining field, recommender systems (RecSys) serve as an indispensable and effective technique in addressing information overload problems and enriching user experience across diverse applications [1, 2, 3], such as e-commerce, job search, and social media platforms. To provide personalized recommendations that accord with user preferences, one of the most representative techniques is collaborative filtering (CF), which aims to capture collaborative knowledge by modeling the history of user-item interactions [4]. For example, as the most classic CF method, Matrix Factorization (MF) [5] decomposes the use-item matrix into two low-rank matrices (i.e., user&item representations) to represent each user and item with a unique ID (namely ID indexing/tokenizing), and calculates matching scores to predict user behaviors and generate recommendations via the inner product between their representations. Due to the superior ability in representation learning on graph-structured data, Graph Neural Networks (GNNs) (e.g., LightGCN [6] and GTN [7]) have been recently employed to significantly enhance CF by capturing higher-order collaborative knowledge on user-item interaction graphs for recommendations in IDs manners. The main idea of the existing methods is to obtain learnable representations (i.e., token embeddings) for discrete user&item IDs (i.e., tokens).
@@
More recently, the rapid development of Large Language Models (LLMs) techniques (e.g., ChatGPT and LLaMA) has showcased notable milestones for revolutionizing natural language processing techniques [8, 9, 10]. Technically, the unprecedented power of LLMs can be attributed to the scaling up of model/parameter size alongside a tremendous amount of training corpus. In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation abilities, along with remarkable generalization capability and in-context learning skills that facilitate LLMs to better generalize to unseen tasks and domains [9]. Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems [3]. For example, ChatRec [11] integrates ChatGPT into conversational recommendations, which enables users to deliver their explicit preferences in natural language. P5 [12] introduces an LLM-based recommendation framework that unifies diverse recommendation tasks by multi-task prompt-based pre-training.
@@ II-C3 $K$-way Encoder and Codebook
$\text{item}~v_{j}$ $\rightarrow\text{tokens: }\{w_{j}^{1},w_{j}^{2},...,w_{j}^{K}\}$ (7)
$\rightarrow\text{tokens' embeddings: }[\mathbf{c}^{1}_{w^{1}_{j}},\mathbf{c}^{2}_{w^{2}_{j}},...,\mathbf{c}^{K}_{w^{K}_{j}}].$ (8)
@@ II-D1 Tokenization&Prompts
Technically, Prompt 1 and Prompt 2 showcase the input prompts without and with item interactions as supporting information, respectively, in which $\left\langle uk\text{-}\cdot\right\rangle$ denotes the out-of-vocabulary (OOV) tokens in the $k_{th}$ sub-codebook. For instance, $\left\langle u2\text{-}21\right\rangle$ represents the $21_{st}$ token in the second sub-codebook for the user_03. This also applies to tokenizing items in LLMs.
@@ II-D2 User Modeling via LLM
where $\mathcal{T}_{t}$ represents the token being generated at the $t_{th}$ position, while $\mathcal{T}_{1:t-1}$ denotes the previously generated tokens from the LLM. In contrast, our user modeling process differs from the previous approach. To be more specific, we take the input $\mathcal{X}_{i}$ and pass it through the LLM backbone, denoted as $\text{LLM4Rec}(\cdot)$ to generate a hidden representation $\mathbf{h}_{i}$ that reflects the model’s comprehension of user $u_{i}$’s preferences for next-items recommendations. Mathematically, we can express this process as follows:
@@
In other words, the final encoded representation $\mathbf{h}_{i}$ can be considered as user $u_{i}$’s generative preferences of the next items for making personalized recommendations. Within our framework, the LLM backbone functions as a powerful query encoder, positioned to excel over traditional deep neural networks in user modeling across several critical dimensions: (a) comprehending personalized user queries for recommendations through diverse prompts with/without user’s historical interactions; (b) interpreting users’ preferences leveraging LLM’s impressive abilities in reasoning; and (c) generating desired outcomes: LLMs can beyond text [26, 27].
@@ II-D3 Generative Retrieval
More specifically, the hidden state $\mathbf{h}_{i}$ from $\text{LLM4Rec}(\cdot)$ will be projected to a latent representation $\mathbf{z}_{i}\in\mathbb{R}^{d}$ through a projection layer $\text{Proj}(\cdot)$, to make the alignment between the LLM-generated representation and item representations learned from GNNs as follows:
@@
To sum up, given the scores of all items in the item base, the proposed TokenRec can easily retrieve top-$K$ items for users rather than directly generating tokens in natural language. This approach offers advantages in terms of efficiency in inference and avoiding hallucinations [29], distinguishing TokenRec from the majority of existing language model-based Recommender Systems that struggle with their time-consuming auto-regressive decoding and beam search processes. Furthermore, unseen items in the fine-tuning stage can be retrieved by only updating the item representations pool $\mathcal{V}$ instead of retraining the entire model. Lastly, this two-tower-like structure [30] shows promise in facilitating seamless alignment between textual information on the query side and collaborative knowledge on the hidden representation side.
@@ II-E1 Training
Figure 3: The TokenRec’s efficiency and generalization capability for new users and items during the inference stage. Rather than retraining the MQ-Tokenizers and LLM backbone, which can be computationally expensive and time-consuming, only the GNN needs to be updated for learning representations for new users and items.
@@ II-E2 Inference
  • Generalizability to New Users and Items. The proposed architecture can provide robust ID tokenization for unseen users and items without fine-tuning the LLM4Rec component. In practice, web platforms often encounter a frequent influx of new products (items) and users. These entities might have a few interactions. For instance, in social media platforms, numerous new users join daily, sharing tweets on diverse topics, following others, and interacting through likes, retweets, and replies. However, these interactions are not encompassed within the training corpus of recommendation models, rendering them as unseen users/items. Maintaining regular updates to the foundational model for this daily influx proves to be a costly endeavor for existing LLM-based RecSys. To address this problem, we propose to leverage external lightweight GNN models to learn corresponding representations of these unseen items/users, so that TokenRec can perform inference on them without requiring retraining of LLM backbone and MQ-Tokenizers. As shown in Figure 3, when new users and items are added to the system, retraining is only required for the collaborative filtering component (e.g., MF and LightGCN) to capture collaborative knowledge in learning representations of these users and items (i.e., updating vector database). In contrast, the MQ-Tokenizers and LLM backbone can remain frozen and perform well, thanks to the masking and $K$-way encoder mechanisms in vector quantization. This addresses the cold-start problem for LLM-based RecSys, eliminates the need for retraining LLM-related components, and therefore saves huge computational resources: compared to the finetuning of LLMs, the training of GNNs is far more efficient [34]. Notably, the capability to adapt to new users and items is demonstrated in Section III-C and Table IV.
@@ III-A2 Baselines
Here, we compare our approach with four representative collaborative filtering methods (i.e., MF, NeuCF, LightGCN, GTN, and LTGNN), three widely-used sequential recommendation methods (i.e., SASRec, BERT4Rec, and S3Rec), and five state-of-the-art LLM-based recommendation methods(i.e., P5, CID, POD, TIGER, and CoLLM).
  • Collaborative Filtering (CF) methods: MF [38] is the most classic CF method, while NeuCF [39] is the very first DNN-based model. LightGCN [6] and GTN [7] are representative CF based on GNNs techniques. LTGNN [34] is the most advanced GNN-based collaborative filtering method.
  • Sequential Recommendations: SASRec [40] is an attention-based sequential recommendation model. BERT4Rec [41] is a bidirectional Transformer-based recommender trained with the BERT-style cloze task. S3Rec [42] is a representative sequential recommendation model trained by self-supervised learning. CoSeRec [43] incorporates contrastive Self-Supervised Learning to sequential recommendation.
@@
  • CID [44] is an indexing approach that considers the co-occurrence matrix of items to design numeric IDs so that items co-occur in user-item interactions will have similar numeric IDs. To be consistent, we employ the P5 model as its LLM backbone.
@@
  • TIGER [46] condenses extensive textual data into a few semantic IDs through residual vector quantization and trains a Transformer-based model using the sequences of semantic IDs for sequential recommendation tasks. A variant is also included in our evaluation, denoted by TIGER-G. It is trained on collaborative IDs instead of semantic IDs. The collaborative IDs are quantized from the collaborative embeddings learned by a GNN method, specifically LightGCN.
  • CoLLM [16] employs GNNs to provide continuous embeddings representing items and users for LLM-based recommendations. In our experiments, CoLLM’s binary classification output (i.e., whether or not a user likes an item) is reformatted to generate item IDs consistent with the output setting of top-$K$ recommendations.
@@ III-A3 Evaluation Metrics
In order to evaluate the quality of recommendation results, we adopt two widely used evaluation metrics: top-$K$ Hit Ratio (HR@$K$) and top-$K$ Normalized Discounted Cumulative Gain (NDCG@KK[20, 34], in which higher values indicate better performance for recommendations. And the average metrics for all users in the test set are reported. In addition, we set the values of $K$ as 10, 20, and 30, among which 20 is the default value for ablation experiments.
@@ III-A4 Hyper-parameter Settings
Our proposed model is implemented based on Hugging Face and PyTorch. The codebook number $K$ (i.e., the subencoder), the token number $L$ at each subcodebook, and the ratio $\rho$ of our masking operation are search in the ranges of $\{1,2,3,4,5\}$, $\{128,256,512,1024\}$, and {0 to 1} in 0.1 increments, respectively. Moreover, during the fine-tuning process of our proposal LLM-based recommendation framework, the ratio of negative sampling $\lambda$ presented in Eq.(20) is fixed at 1:1, in which we randomly select an un-interacted item from the whole item base as the negative sample for each positive sample. The margin $\gamma$ in Eq.(20) is set to 0 to 0.2. We optimize the MQ-Tokenizers and the LLM backbone with AdamW [47] in a mini-batch manner, with a batch size of 128 and a maximum of 100 training epochs. Note that the high-order collaborative representations for users and items are obtained from a representative collaborative filtering method, namely LightGCN [6]. For prompting, we design 11 templates for TokenRec: 10 of them are set to be seen prompts, and the remaining one is evaluated as an unseen prompt. For a fair comparison, we employ a widely-used lightweight LLM, i.e., T5-small [37], for TokenRec and all LLM-based baselines. The other default hyper-parameters for baseline methods are set as suggested by the corresponding papers. All the experiments are conducted on a single NVIDIA A800 GPU (80 GB).
@@ III-B Performance Comparison of Recommender Systems
  • As an earlier non-trivial indexing method to capture hidden knowledge in co-occurrence frequency, Collaborative Indexing (CID) outperforms Random Indexing (RID) and Sequential Indexing (SID) given the same setting of P5. These observations suggest the potential of integrating collaborative knowledge for item&user tokenization/indexing. However, the P5 variants and POD are inferior to the existing GNN-based collaborative filtering (i.e., LightGCN, GTN, and LTGNN), implying their inability to capture collaborative information using LLMs. In constrast, CoLLM achieves superior results, benefiting from its incorporation of collaborative embeddings learned from these GNN methods.
  • Thanks to its use of residual vector quantization and semantic IDs for item indexing, TIGER outperforms typical LLM-based recommendation methods, such as P5-RID and P5-SID. Moreover, the variant TIGER-G demonstrates improved performance in many cases compared to TIGER, further highlighting the benefits of incorporating collaborative knowledge in LLM-based recommendation tasks.
  • The GNN-based collaborative filtering methods perform relatively better than the traditional CF methods (i.e., MF and NCF) and representative sequential recommendation methods (i.e., BERT4Rec, SASRec, S3Rec, and CoSeRec). The results demonstrate the effectiveness of GNNs in capturing collaborative signals via high-order connectivity.
@@ III-C Generalizability Evaluation
For this analysis, we consider the Beauty and LastFM datasets and exclude the 5% of users (ref as unseen users) with the least interaction history from the training data split to simulate newly added users. Consequently, all items interacted with by these users are also excluded from the training set. The performance of TokenRec is compared to other LLM-based recommendation models, i.e., P5, POD, CID, TIGER, and CoLLM. To ensure the absence of data leakage concerning unseen users, only the training split is utilized for establishing these models, while it is allowed to update the vector database to provide collaborative representations of unseen users and items (as shown in Figure 3), which consumes far less computational resources than updating the LLM backbone. From the recommendation results presented in Table IV, several key observations can be made:
@@
  • In comparison, TokenRec outperforms the aforementioned methods for not only the training users but also the unseen users in both datasets. For instance, in the Amazon-Beauty dataset, the performance of TokenRec decreases by only 7% on average, demonstrating the strong generalization capability of TokenRec for newly added users. Such a superiority can be attributed to our MQ-Tokenizer for robust ID tokenization and the generative retrieval paradigm for flexible recommendation generation.
  • Additionally, TIGER exhibits strong generalization capability for unseen cases, with only a 6.10% average decrease observed on HR@20. This performance can be attributed to the use of semantic IDs, which incorporate item-side textual information as additional knowledge—a feature not present in other compared methods. This approach holds potential for addressing cold-start scenarios by incorporating item-side information effectively.